Insourcing incurs a periodic fixed cost of $20,000 and a $4.50 variable cost. Outsourcing incurs a periodic cost of $45,000 and a $2.50 variable cost. Over what ranges of demand is each option best?

A.Outsourcing is optimal for demand > 25,000 units. Insourcing is optimal for demand < 25,000 units

B.Outsourcing is optimal for demand < 12,500 units. Insourcing is optimal for demand > 12,500 units

C.Insourcing is optimal for demand > 12,500 units. Outsourcing is optimal for demand > 12,500 units

D.Insourcing is optimal for demand > 25,000 units. Outsourcing is optimal for demand < 25,000 units

E.Outsourcing is optimal for demand > 12,500 units. Insourcing is optimal for demand < 12,500 units
Answer #1

Let x be the point at which insourcing cost is equal to the outsourcing cost

Insourcing cost = 20000 + 4.5x

Outsourcing cost = 45000 + 2.5x

=> 20000 + 4.5x = 45000 + 2.5x

=> x = 25000/2 = 12500

Hence, Insourcing is optimal for demand < 12500 and outsourcing is optimal for demand > 12500

Hence, option (e)